• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 07:14:50 on October 2,1999

    Entry number 27440

    keyword=config change

    At the previous e-P setting, which finished 6:15, we took 3.5 M events.
    Now we have E=2.5, P_e=1.9014, P_h=1.2160. Measured angles are
    E=28.121, H=47.469. H-dipole settled in 45 minutes. MCC informs us
    they have a cold compressor crash. No beam expected or awhile.